OpenStudy (anonymous):

one brand of cheddar cheese sells for $0.29 per ounce. Tanya bought a block of the cheese for $2.61. What equation could be used to determine the weight of the cheese in ounces?

OpenStudy (anonymous):

Hacklanding i think that, one block of cheese's weight is 1 pound and all you need to figure out is the ounces of the pound and add the money and see what equation you used.

OpenStudy (anonymous):

.29 is 1 pound 2.61/0.29 =9 pound

OpenStudy (anonymous):

my choices are w/2.61=0.29 w/0.29=2.61 2.61w=0.29 0.29w=2.61

OpenStudy (anonymous):

9 is the answers

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

0.29w/0.29=2.61/0.29 w=2.61/0.29 w=9

OpenStudy (anonymous):

np